# Enhancing Large Vision Language Models with Self-Training on Image Comprehension

### v1

@@ 5.2 Main Results
Model ScienceQA TextVQA ChartQA LLaVA-Bench MMBench MM-Vet MathVista
InstructBLIP (7B) 60.5 50.1 60.9 36.0 26.2 25.3
mPLUG-OWL2 (7B) 64.5 54.3 59.9 64.5 36.2 22.2
LLaVA-v1.5 (7B) 66.8 58.2 6.32 65.4 64.3 31.1 25.1
w/ POVID 68.8 68.7 64.9 31.8
w/ STIC 69.5 61.4 6.64 68.9 65.3 32.6 27.2
@@
We present our main results in Table 1 and detail the benchmark performances of STIC (LLaVA-v1.6 7B) on MMBench and MM-Vet in Figure 4. In Appendix A, we present detailed results for MMBench in Table B and MM-Vet in Table B. Our results show a consistent and significant improvement of STIC over the original models (LLaVA-v1.5 and LLaVA-v1.6) across all seven datasets. This improvement is achieved using only self-constructed preference data and a small portion of the model’s SFT dataset, which had already been used for fine-tuning the original model.
On average, STIC improves LLaVA-v1.5 by $1.7\%$, increasing from $45.3\%$ to $47.0\%$. and LLaVA-v1.6 by a notable score of $4.0\%$, increasing from $54.7\%$ to $58.7\%$. The improvement is comprehensive, as detailed in Tables B and B, where STIC consistently enhances performance across all evaluation tasks and targets. Moreover, while STIC improves both LLaVA-v1.5 and LLaVA-v1.6, a more significant improvement is observed in the more advanced model, LLaVA-v1.6. This trend suggests that the extent of self-improvement could be correlated with the model’s inherent capabilities.
@@ 6 Ablation Studies and Discussions
In this section, we outline the differences of our method with current alignment fine-tuning methods focusing on the vision-language setting. Furthermore, we conduct ablation studies on the key components of STIC to demonstrate their importance and effectiveness. Additionally, we examine the image distribution of our self-training data (MSCOCO) alongside the image distributions of benchmark datasets, revealing a positive correlation between performance gains and similarity in image distributions.
Figure 5: Data comparison.
Discussion with POVID.
Effectiveness of describe-and-respond (DaR) prompting.
We assess the significance of the fine-tuning process in STIC by comparing it to the approach of directly allowing the base LVLM to describe an image and then respond to a query with a self-augmented prompt, which we refer to as the describe-and-respond (DaR) prompting method. As indicated in Table 2, applying DaR to the base LVLM yields mixed results, with performance improvements on some datasets and degradation on others, resulting in an overall average drop of $2.2\%$. In contrast, when DaR is combined with the fine-tuning process of STIC, it leads to a further average enhancement of $1.1\%$ and a notable $2.8\%$ on ScienceQA. This demonstrates the synergistic effect of DaR and the fine-tuning process in STIC. Additionally, it is worth noting that STIC achieves a substantial average improvement of $2.9\%$ even without the DaR prompting method, compared to the base LVLM.
@@
Method DaR ScienceQA TextVQA ChartQA LLaVA-Bench MMBench MM-Vet MathVista Average
Original 68.9 60.3 36.4 77.3 63.7 42.2 34.6 54.8
69.9 56.6 34.6 78.5 50.7 42.3 34.7 52.5
w/ STIC 72.5 63.4 39.3 78.4 68.7 45.7 35.2 57.6
@@
The role of dispreferred samples in STIC.
To understand the importance of dispreferred samples in STIC, we conduct an ablation study using llava-v1.6-mistral-7b as the base LVLM. We remove the negative examples from the preference data and only use the positive samples for supervised fine-tuning (SFT), effectively creating an SFT version of STIC. Table 3 shows that omitting the dispreferred samples evev leads to a performance drop of $0.6\%$ on LLaVA-Bench, while failing to provide equally significant improvement as STIC with preference data. This highlights the crucial role of negative examples in aligning preferences and enabling the model to distinguish between high-quality and low-quality responses. By leveraging both positive and negative examples, STIC effectively improves the model’s performance and generates more preferred outputs.
@@
Model ScienceQA TextVQA LLaVA-Bench
@@
Figure 6: Progression of stages in STIC.
Progression of stages.
In Figure 6, we illustrate the sequential improvement in performance of STIC on ScienceQA. While stage 1 focuses exclusively on enhancing the perception capabilities of the LVLM, it still notably improves performance on downstream VQA tasks. Building on the improved image comprehension achieved in stage 1, stage 2 introduces an enhanced reasoning process that utilizes the model’s self-generated image descriptions and results in an even more significant gain. This enhancement further enables the model to self-augment its prompts with Describe and Respond (DaR), resulting in total the substantial performance gains of $6.4\%$ observed.
@@
Scaling law of STIC.
We explore the scaling law of STIC by expanding the preference data in Stage 1. Using the LLaVA-Bench benchmark as a case study, we scale up the preference data from $6k$ to $12k$ MSCOCO images. As depicted in Figure 7, there is an obvious gain on the LLaVA-Bench from $1.9\%$ to $3.1\%$. This finding demonstrates that STIC can effectively leverage larger amounts of unlabeled image data and presents a cost-effective solution to the challenge of acquiring high-quality vision-language data.
Figure 8: t-SNE visualization of images from MSCOCO and four benchmarks, each sampling $1k$ images.
Correlation between image distribution and performance gains.
To gain further insight into the effectiveness of STIC across different benchmarks, we conducted a t-SNE visualization analysis comparing the image distributions of MSCOCO, which we used for preference data construction, with those of four benchmarks: ScienceQA, TextVQA, MathVista, and ChartQA (Figure 8). Our analysis revealed a general trend: the greater the overlap between the MSCOCO image distribution and that of a benchmark, the higher the performance gain achieved by STIC on that benchmark. This observation held true for ScienceQA and TextVQA, which exhibited substantial distributional overlap with MSCOCO and yielded the highest performance gains of $6.4\%$ and $4.9\%$, respectively. Conversely, MathVista, with its diverse image types and limited overlap with MSCOCO, saw a more modest gain of $2.4\%$. Interestingly, ChartQA was an outlier, achieving a high gain of $5.1\%$ despite minimal overlap with MSCOCO, suggesting that the improved image comprehension from STIC played a fundamental role in understanding and reasoning about the charts. Detailed per-benchmark visualizations and discussions are provided in Appendix B.2.
@@ 7 Conclusion
The promising results demonstrated by STIC in enhancing the capabilities of 7B LVLMs suggest its potential applicability to larger models, such as those with 13B, 40B, and 100B parameters, if computational resources permit. Additionally, it would be beneficial to investigate the impact of the image distribution used in self-training on STIC, aiming to further refine its effectiveness with curated image datasets. Lastly, an examination of the effects of various image corruptions and “bad” prompts on STIC could provide valuable insights into the effective generation of dispreferred samples.
Appendix A Experimental Details
@@
Appendix B Experimental Results
@@ B.1 Example Outputs
Appendix C Limitations and Future Work
@@
Appendix D Broader Impacts

### v2

@@ 5.2 Main Results
{tblr}
colspec = l c c c c c c c, row1 = bg=gray!25, row2,4,6 = bg=gray!10, row8 = bg=LightCyan, Model ScienceQA TextVQA ChartQA LLaVA-Bench MMBench MM-Vet MathVista
InstructBLIP (7B) 60.5 50.1 – 60.9 36.0 26.2 25.3
mPLUG-OWL2 (7B) 64.5 54.3 – 59.9 64.5 36.2 22.2
LLaVA-v1.5 (7B) 66.8 58.2 6.3 65.4 64.3 31.1 25.1
w/ POVID 68.8 – – 68.7 64.9 31.8 –
w/ STIC 69.5 61.4 6.6 68.9 65.3 32.6 27.2
@@
We present our main results in Table 5.2 and detail the benchmark performances of STIC (LLaVA-v1.6 7B) on MMBench and MM-Vet in Figure 4. In Appendix B, we present detailed results for MMBench in Table C and MM-Vet in Table C. Our results show a consistent and significant improvement of STIC over the original models (LLaVA-v1.5 and LLaVA-v1.6) across all seven datasets. This improvement is achieved using only self-constructed preference data and a small portion of the model’s SFT dataset, which had already been used for fine-tuning the original model.
On average, STIC improves LLaVA-v1.5 by $1.7\%$, increasing from $45.3\%$ to $47.0\%$, and LLaVA-v1.6 by a notable score of $4.0\%$, increasing from $54.7\%$ to $58.7\%$. The improvement is comprehensive, as detailed in Tables C and C, where STIC consistently enhances performance across all evaluation tasks and targets. Moreover, while STIC improves both LLaVA-v1.5 and LLaVA-v1.6, a more significant improvement is observed in the more advanced model, LLaVA-v1.6. This trend suggests that the extent of self-improvement could be correlated with the model’s inherent capabilities.
@@ 6 Ablation Studies and Discussions
In this section, we conduct ablation studies on the key components of STIC to demonstrate their importance and effectiveness. Additionally, we examine the image distribution of our self-training data (MSCOCO) alongside the image distributions of benchmark datasets, revealing a positive correlation between performance gains and similarity in image distributions.
Effectiveness of describe-and-respond (DaR) prompting. We assess the significance of the fine-tuning process in STIC by comparing it to the approach of directly allowing the base LVLM to describe an image and then respond to a query with a self-augmented prompt, which we refer to as the describe-and-respond (DaR) prompting method. As indicated in Table 6, applying DaR to the base LVLM yields mixed results, with performance improvements on some datasets and degradation on others, resulting in an overall average drop of $2.3\%$. In contrast, when DaR is combined with the fine-tuning process of STIC, it leads to a further average enhancement of $1.1\%$ and a notable $2.8\%$ on ScienceQA. This demonstrates the synergistic effect of DaR and the fine-tuning process in STIC. Additionally, it is worth noting that STIC achieves a substantial average improvement of $2.8\%$ even without the DaR prompting method, compared to the base LVLM.
@@
{tblr}
colspec = l c | c c c c c c c | c, row1 = bg=gray!25, row4-5 = bg=gray!10, Method DaR ScienceQA TextVQA ChartQA LLaVA-Bench MMBench MM-Vet MathVista Average
[…]
✓ 69.9 56.6 34.6 78.5 50.7 42.3 34.7 52.5
[…]
@@
Figure 5: Progression of stages in STIC.
Progression of stages. In Figure 5, we illustrate the sequential improvement in performance of STIC on ScienceQA. While stage 1 focuses exclusively on enhancing the perception capabilities of the LVLM, it still notably improves performance on downstream VQA tasks. Building on the improved image comprehension achieved in stage 1, stage 2 introduces an enhanced reasoning process that utilizes the model’s self-generated image descriptions and results in an even more significant gain. This enhancement further enables the model to self-augment its prompts with Describe and Respond (DaR), resulting in total the substantial performance gains of $6.4\%$ observed.
The role of dispreferred samples in STIC. To understand the importance of dispreferred samples in STIC, we conduct an ablation study using llava-v1.6-mistral-7b as the base LVLM. We remove the negative examples from the preference data and only use the positive samples for supervised fine-tuning (SFT), effectively creating an SFT version of STIC. Table 6 shows that omitting the dispreferred samples even leads to a performance drop of $0.6\%$ on LLaVA-Bench, while failing to provide equally significant improvement as STIC with preference data. This highlights the crucial role of negative examples in aligning preferences and enabling the model to distinguish between high-quality and low-quality responses. By leveraging both positive and negative examples, STIC effectively improves the model’s performance and generates more preferred outputs.
@@
{tblr}
colspec = l c c c , row1 = bg=gray!25, row3 = bg=gray!10, Model ScienceQA TextVQA LLaVA-Bench
@@
Scaling law of STIC. We explore the scaling law of STIC by expanding the preference data in Stage 1. Using the LLaVA-Bench benchmark as a case study, we scale up the preference data from $6k$ to $12k$ MSCOCO images. As depicted in Figure 6, there is an obvious gain on the LLaVA-Bench from $1.9\%$ to $3.1\%$. This finding demonstrates that STIC can effectively leverage larger amounts of unlabeled image data and presents a cost-effective solution to the challenge of acquiring high-quality vision-language data.
Figure 7: t-SNE visualization of images from MSCOCO and four benchmarks, each sampling $1k$.
Correlation between image distribution and performance gains. To gain further insight into the effectiveness of STIC across different benchmarks, we conducted a t-SNE visualization analysis comparing the image distributions of MSCOCO, which we used for preference data construction, with those of four benchmarks: ScienceQA, TextVQA, MathVista, and ChartQA (Figure 7). Our analysis revealed a general trend: the greater the overlap between the MSCOCO image distribution and that of a benchmark, the higher the performance gain achieved by STIC on that benchmark. This observation held true for ScienceQA and TextVQA, which exhibited substantial distributional overlap with MSCOCO and yielded the highest performance gains of $6.4\%$ and $4.9\%$, respectively. Conversely, MathVista, with its diverse image types and limited overlap with MSCOCO, saw a more modest gain of $2.4\%$. Interestingly, ChartQA was an outlier, achieving a high gain of $5.1\%$ despite minimal overlap with MSCOCO, suggesting that the improved image comprehension from STIC played a fundamental role in understanding and reasoning about the charts. Detailed per-benchmark visualizations and discussions are provided in Appendix C.2.
[…]
Table 4: Performance of STIC on different stage-1 training images compared with the original LVLM model LLaVA-v.16 (Mistral 7B) across vision-language reasoning benchmarks.
{tblr}
[…]
Complex Conv Detail All Rec Ocr Know Gen Spat Math Total All
LLaVA-v1.6 (7B) - 87.4 61.3 77.8 77.3 43.1 40.6 29.6 32.5 44.7 15.4 42.2 63.7
w/ STIC COCO 89.1 63.7 79.5 79.2 45.7 42.5 30.4 34.9 45.1 22.7 45.0 67.8
w/ STIC VFLAN 92.8 68.4 77.9 81.9 45.7 43.0 31.0 36.2 45.1 26.5 45.1 68.3
Scalability. To explore STIC’s applicability to models with higher representation capacity, we conducted supplementary experiments using LLaVA-v1.6 (Vicuna-13B). Table 6 shows the detailed experiment results. We used the same images for STIC fine-tuning as in our experiments for LLaVA-v1.6 (Mistral-7B) to ensure fairness and the same set of hyperparameters. The improvements observed with LLaVA-v1.6 (Vicuna-13B) demonstrate that STIC is not only effective with smaller models but also scales well with larger or more capable LVLMs.
Table 5: Performance of STIC compared with the original LVLM model LLaVA-v1.6 (Vicuna 13B) across vision-language reasoning tasks. Image data used for 13B model remain the same as what we used for the 7B model.
{tblr}
[…]
Complex Conv Detail All Rec Ocr Know Gen Spat Math Total All
LLaVA-v1.6 (7B) 87.4 61.3 77.8 77.3 43.1 40.6 29.6 32.5 44.7 15.4 42.2 63.7
LLaVA-v1.6 (13B) 94.0 73.8 78.7 84.5 52.2 47.1 38.8 45.2 42.7 26.9 48.9 70.6
w/ STIC 93.5 78.1(+4.3) 79.4 85.6(+1.1) 54.5 48.0 42.3(+3.5) 49.4(+4.2) 42.0 23.1 50.5(+1.6) 72.3(+1.7)
Qualitative example. In Figure 8, we show an example output of STIC. Despite the task being focused on mathematical reasoning, STIC enhanced the model’s response by improving its image comprehension capabilities. While the original model merely identified one of the recognized numbers in the image as the final answer, the STIC fine-tuned model was able to interpret the meaning of each number, describe them accurately, and perform reasoning based on this understanding.
@@ 7 Conclusion
The promising results demonstrated by STIC in enhancing the capabilities of 7B LVLMs suggest its potential applicability to larger models, such as those with 40B, and 100B parameters, if computational resources permit.
Contribution Statement
The project began in January 2024 with Deng as part of Gu’s group, collaborating with Lu from Chang’s group. This project was in the line of research on self-training using synthetic data developed by Gu and Deng. Deng proposed the initial idea of this project, which was further developed with Lu. In February 2024, Yin joined the project. Both Gu and Chang offered early feedback to help shape the project. In March, Chang met with Deng, Lu, and Yin to discuss the initial experiment results and finalize the research plan. Chang continued supervising the project through one-on-one meetings with Lu. In April, Deng changed advisors and moved to Wang’s lab, inviting Hu, Shen, Wang, and Zou to join the project. Hu and Shen contributed algorithmic improvements, while Wang and Zou provided detailed feedback on the experimental design and ablation studies. Deng and Lu primarily conducted the experiments and drafted the paper.
Acknowledgments
We sincerely thank the anonymous reviewers for their helpful comments. Pan is supported by Bloomberg Data Science Ph.D. Fellowship, Qualcomm Innovation Fellowship and UCLA Dissertation Year Fellowship. Fan and Kai-Wei are supported by ONR grant N00014-23-1-2780, OptumLabs, and CISCO gift grants. Wei is supported by DARPA HR0011-24-9-0370, NSF 2200274, 2106859, 2312501, and NIH U54HG012517, U24DK097771. This paper is partially supported by Chan Zuckerberg Initiative. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
Appendix A Additional Related Work
Self-improving language models. High-quality data, including human-crafted and advanced AI generated content, has been demonstrated to significantly enhance the performance of LLMs on various tasks (Josifoski et al., 2023; Taori et al., 2023; Chiang et al., 2023a; Li et al., 2023c). Although, acquiring such high-quality data is often prohibitively expensive. To circumvent the costs associated with obtaining human-annotated or expertly curated data, researchers have shifted their focus to leveraging data generated by the target model itself, exploring ways of self-improvement (Chen et al., 2024; Yuan et al., 2024; Fränken et al., 2024; Rosset et al., 2024). Recent studies have also emphasized the rephrasing capabilities of LLMs, which either enhance their own response quality (Deng et al., 2023; Prasad et al., 2023) or augment synthetic data for self-supervised fine-tuning (Kim et al., 2023). To the best of our knowledge, our work is the first to explore the potential for self-improvement in LVLMs, specifically focusing on the vision modality and emphasizing the self-improvement of image comprehension capabilities.
Appendix B Experimental Details
@@
The prompts used as instructions in DPO are listed below:
  • “Illustrate the details of the picture.”
  • “Summarize the visual content presented.”
  • “Explain what is depicted in the photograph.”
  • “Outline the key elements captured in the image.”
  • “Detail the composition and subjects within the frame.”
  • “Convey the atmosphere and mood represented in the snapshot.”
  • “Interpret the scene shown in the image.”
  • “Identify and describe the main focal points in the visual.”
@@
Describe and Respond.
[…]
Model: <image description>
[…]
Model: <response>
@@
Appendix C Experimental Results
@@ C.1 Example Outputs
C.3 Discussion with POVID.
Figure 15: Data comparison.
C.4 Investigation of Prompt Quality
Table 9 presents an additional experiments using DaR to demonstrate prompt quality. We compared normal prompts from our main paper (e.g., “Illustrate the details of the picture.”) with the hallucination prompts and well-curated prompts used for DPO pair generation. The results show an expected discrepancy in QA performance: hallucination prompts significantly decreased performance, while well-curated prompts maintained a decent performance. We also included results based on a prompt proposed by Llama-3 8B and filtered using the same restrictions. The performance difference between GPT-4 and Llama-3 8B prompts underscores the quality of GPT-4’s proposals.
Table 9: Test performance of llava-v1.6-mistral-7b using various prompts with DaR. We evaluate prompt quality using DaR as a prompting method. DaR=None represents the original LVLM model’s performance. Normal prompt refers to the simple prompt we used for DaR in our paper. GPT-4’s well-curated prompt refers to the prompt we used for preferred response generation, and we include Mistral 7B’s curated prompt for additional comparison.
Model DaR LLaVA-Bench MM-Vet MMBench
LLaVA-v1.6 (7B) None 77.3 42.2 63.7
Normal Prompt 78.5(+1.2) 42.3(+0.1) 50.7(-13.0)
Hallu Prompt 73.7(-3.6) 40.5(-1.7) 40.7(-23.0)
Well-curated (Llama-3 8B) 77.2(+0.1) 40.0(-2.2) 60.1(-3.6)
Well-curated (GPT-4) 79.1(+2.1) 42.9(+0.7) 60.9(-2.8)
Appendix D Limitations and Future Work
@@
Appendix E Broader Impacts